William is planning to create a rectangular mosaic which measures 65 cm by 95 cm. The mosaic will be covered completely with square pieces of colored glass. William has decided that he will purchase only one size of glass squares, and he does not plan to cut any of the pieces. If the art supply store only sells the glass squares in whole-number side lengths (measured in centimeters), find the smallest number of squares which William could use for his mosaic.
Answer by ikleyn(52778)

.
Greatest common factor (GCF) of the numbers 65 and 95 is 5:

65 = 5*13,
95 = 5*19.

So, the elementary (atomary) squares shoud be 5x5 and their number should be 13*19 = 247.